Find the exact value of the trigonometric function at the given real number.
(a) cos(− π/3) =
(b) sec(− π/3) =
(c) tan(−π/3) =
Please explain
Thank you

Answer by ewatrrr(24785)

(a) cos(− π/3) = cos(π/3) = 1/2
(b) sec(− π/3) = 1/cos(π/3) = 2/1
(c) tan(−π/3) = - tan(π/3) = -√3
Note: tanx = sinx/cosx
.....See below (cosx, sinx) Summary Unit Circle
.